# supplement

### v1

@@ 5.2 Example 2: A stochastic differential equation example
We consider parameter estimation of a high dimensional non-linear example, the Van der Pol oscillator model (Kandepu et al., 2008; Särkkä et al., 2015). The model is described by the following second-order non-linear ODE:
@@
Here $\omega_{c}$ represents the angular velocity of the force process, $\sigma_{n}$ represents the strength of the noise process driving the $n$th harmonic, and $\epsilon_{n}(t)$ represents a white noise process. In this section, we set $N=2$. For simplicity, we assume that $\sigma_{n}=\sigma_{c}$ for all $n=1,2$. Consequently, the parameter vector $\boldsymbol{\theta}=(\epsilon,\mu,\sigma,\omega_{c},\sigma_{c})$ is five-dimensional, and the state variable $\mathbf{x}=(x,\dot{x},c_{1},\dot{c}_{1},c_{2},\dot{c}_{2})$ has $6$ components. The full SDE model can be represented as:
@@
Here, $dW_{n}(t)$ represents the differential of a standard Wiener process. The measurements are obtained by adding noise to the state of the Van der Pol oscillator:
@@
The observed data $\boldsymbol{y}$ was simulated from the system using the parameter values $\boldsymbol{\theta}=(1,1/2,1/10,$ $\pi/5,1/100)$ over the time interval $t\in[0,40]$, with a sampling period of $\Delta t=1$. Since ${\boldsymbol{\theta}}\geq 0$, we take Gamma distributions as prior distributions $\pi({\boldsymbol{\theta}})$: $\epsilon\sim Gamma(5,1)$, $\mu\sim Gamma(3,5)$, $\sigma\sim Gamma(5,15)$, $\omega_{c}\sim Gamma(5,10)$, $\sigma_{c}\sim Gamma(2,15)$. Here, we choose the kernel function $K_{\varepsilon}(\cdot,\boldsymbol{y})=N(\cdot;\boldsymbol{y},\varepsilon^{2})$ with and ABC threshold $\varepsilon=0.15$.
We obtain the true posterior by running 5 ABC-MCMC chains with $1,000,000$ iteration, and discard the first 50,000 iterations as the burn-in. We use “density” function in R to estimate the 1D marginal density with 512 points on the posterior region, specifically is $[0,12]\times[0,1.5]\times[0,1]\times[0,1.5]\times[0,0.5]$. The KL divergence is calculated as follows:
@@
where $\pi_{\varepsilon}(\theta_{i})$ is the true ABC posterior marginal density of $\theta_{i}$, $\hat{\pi}_{\varepsilon}(\theta_{i})$ is the estimated ABC posterior marginal density of $\theta_{i}$, and $l_{i}$ is the length of the interval for the posterior estimate. Here, we examine the robustness of our sequential optimization algorithm for selecting $\gamma$ by implementing GL-ABC-MCMC with two different ABC-i-SIR proposal distributions. The batch size of ABC-i-SIR is set to $N_{b}=20$. In the first scenario, we use the prior distribution as the importance proposal, denoted as GLMCMC1. In the second scenario, we use a flatter density as the proposal distribution, denoted as GLMCMC2. We use cESJD to select the global frequency $\gamma$. The detailed setups and results are shown in Supplementary LABEL:supp:sec-SDE. We select $\gamma=1$ for GLMCMC1 and $\gamma=0.8$ for GLMCMC2 based on the cESJD. The result implies that the global move of GLMCMC1 performs sufficiently well, whereas GLMCMC2 requires a combination of global and local moves to maximize the efficiency of the algorithm. The proportion of global moves of GLMCMC1 is higher than GLMCMC2 since the importance proposal of GLMCMC1 is closer to the target compared with the GLMCMC2. The result also indicates the effectiveness of the cESJD criterion and the robustness of the GL-ABC-MCMC algorithm with respect to the hyperparameter $\gamma$.
Figure 7: Comparasion of three ABC-MCMC methods (ABC-MCMC, ABC-i-SIR with batch size 20, and GL-ABC-MCMC with batch size 20 and global frequency $0.1$ and $0.3$). Error bars denote $95\%$ confidence intervals for $D_{KL}$ across 10 replicates. First row: $\epsilon$, $\mu$, $\sigma$; Second row: $\omega_{c}$, $\sigma_{c}$.
We compute the KL divergence of different methods at iteration numbers $10^{5}$, $2\cdot 10^{5}$, $\ldots$, $10^{6}$. We take the time executing one MCMC iteration as the baseline. Figure 7 shows the KL divergence of different methods varying with time cost. The methods proposed in this article (i.e., GL-ABC-MCMC) converge faster than ordinary MCMC, and admit smaller variance. The pairwise marginals of the parameters are shown in Supplementary LABEL:supp:sec-SDE, Figure LABEL:fig:_SDE_2d.
@@ 6 Conclusion
In this article, we consider the task of Bayesian inference for models with intractable likelihood function. Markov chain Monte Carlo methods are often combined with ABC to accelerate the likelihood-free inference. MCMC with local moves often struggles to explore posterior distribution with complex surface. Hence, we design efficient global and local moves to explore the challenge posterior surfaces in the likelihood-free framework.
A likelihood free version of i-SIR is developed to serve as global proposals of ABC-MCMC algorithms, and we propose using sequential optimization algorithm to select the hyper-parameters based on a unit cost version of ESJD. We also prove the V-geometric ergodicity of the ABC Global-Local Markov kernel, which indicates that under certain conditions, the mixing rate of the proposed Markov Kernel is significantly better than the corresponding mixing rate of the local kernel.
Two adaptive schemes are designed to improve the efficiency of our proposed ABC-MCMC. In our first adaptive scheme, we incorporate normalizing flow into ABC-i-SIR to train the proposal distribution. In the second adaptive scheme, we design an ABC version of MALA algorithm, in which the gradient of log-likelihood function is evaluated using the common random numbers technique. Numerical experiments demonstrate that our algorithm outperforms existing methods in synthetic and real application models.
There are several lines of extensions and improvements for future work. The ABC tolerance $\epsilon$ is pre-specified before running our ABC-MCMC. Del Moral et al. (2012) provide a framework for adaptively selecting the sequence of tolerance for ABC within the sequential Monte Carlo (SMC). Our first line of future work is to combine our Global-Local moves with ABC-SMC and investigate adaptive schemes to tune the parameters in the moves. The computation of MH acceptance probability $\alpha$ involves generating synthetic data from the simulation model, which may cause a heavy computational burden for complex problems. Our second line of future work is to accelerate the ABC-MCMC inference speed by using approximations (e.g. Cao et al. (2024); Rodrigues et al. (2020)). A large number of simulations are required to evaluate the gradient function of approximate Bayesian methods for high-dimensional parameter spaces. Our third line of future research is to generate control variates for variance reduction and determine theoptimal simulation size for gradient approximation.
Supplementary Materials
S.1:
@@
Acknowledgements
This work was supported by the National Natural Science Foundation of China (12131001 and 12101333), the startup fund of ShanghaiTech University, the Fundamental Research Funds for the Central Universities, LPMC, and KLMDASR. The authorship is listed in alphabetic order.

### v2

@@ 6.2 Example 2: A stochastic differential equation example
We consider parameter estimation of a high dimensional non-linear example, the Van der Pol oscillator model (kandepu2008applying; sarkka2015posterior). The model is described by the following second-order non-linear ODE:
@@
Here $\omega_{c}$ represents the angular velocity of the force process, $\sigma_{n}$ represents the strength of the noise process driving the $n$th harmonic, and $\epsilon_{n}(t)$ represents a white noise process. In this section, we set $N=2$. For simplicity, we assume that $\sigma_{n}=\sigma_{c}$ for all $n=1,2$. Consequently, the parameter vector $\boldsymbol{\theta}=(\epsilon,\mu,\sigma,\omega_{c},\sigma_{c})$ is five-dimensional, and the state variable $\boldsymbol{x}=(x,\dot{x},c_{1},\dot{c}_{1},c_{2},\dot{c}_{2})$ has $6$ components. The full SDE model can be represented as:
@@
Here $dW_{n}(t)$ represents the differential of a standard Wiener process. The measurements are obtained by adding noise to the state of the Van der Pol oscillator:
@@
The observed data $\boldsymbol{y}$ was simulated from the system using the parameter values $\boldsymbol{\theta}=(1,1/2,1/10,$ $\pi/5,1/100)$ over the time interval $t\in[0,40]$, with a sampling period of $\Delta t=1$. Since ${\boldsymbol{\theta}}\geq 0$, we take Gamma distributions as prior distributions $\pi({\boldsymbol{\theta}})$: $\epsilon\sim Gamma(5,1)$, $\mu\sim Gamma(3,5)$, $\sigma\sim Gamma(5,15)$, $\omega_{c}\sim Gamma(5,10)$, $\sigma_{c}\sim Gamma(2,15)$. Here, we choose the kernel function $K_{\varepsilon}(\cdot,\boldsymbol{y})=N(\cdot;\boldsymbol{y},\varepsilon^{2})$ with ABC threshold $\varepsilon=0.2$.
We compare the efficiency of ABC-MCMC, ABC-i-SIR, and GL-ABC-MCMC. The importance proposal distribution of ABC-i-SIR is the prior distribution and the batch size is 20. We choose the global frequency $\gamma$ based on cESJD.
Table 4: cESJD values under different global frequency parameter.
$\gamma$ 0 0.2 0.4 0.6 0.8 1.0
cESJD 0.0024 0.0060 0.0096 0.0108 0.0150 0.0204
Here, we consider the simulation time cost and choose the optimal global frequency $\gamma=1$ from candidate set shown in Table 4. Additionally, we select a ‘worst’ $\gamma=0$ and a ‘sub-worst’ one $\gamma=0.2$ as references. We obtain the true posterior by running 5 ABC-MCMC chains with $1,000,000$ iteration, and discard the first 50,000 iterations as the burn-in. Here, we use “density” function in R to estimate the 1D marginal density with 512 points on the posterior region, that is $0,12]\times[0,1.5]\times[0,1]\times[0,1.5]\times[0,0.5]$. And the KL divergence is calculated by
@@
where $\pi_{\varepsilon}(\theta_{i})$ is the true ABC posterior marginal density of $\theta_{i}$, $\hat{\pi}_{\varepsilon}(\theta_{i})$ is the estimated ABC posterior marginal density of $\theta_{i}$, and $l_{i}$ is the length of the interval for the posterior estimate. Here we calculate KL divergence of different methods at iteration numbers $10^{5}$, $2\cdot 10^{5}$, $\ldots$, $10^{6}$. We take the time executing one MCMC iteration as the baseline. Figure 10 shows the KL divergence of different methods varying with time cost. The methods proposed in this article (i.e., GL-ABC-MCMC and ABC-i-SIR) converge faster than ordinary MCMC, and admit smaller variance. Figure 11 shows the one-dimensional marginal distribution of the parameters. Figure 12 shows the results correspond to the first 20,000 iterations out of 500,000 total iterations. These results reveal that ABC-MCMC tends to get trapped in local modes, resulting in insufficient coverage of the tails (tail deficiency). Moreover, it may spend excessive time dwelling in low-probability regions, which adversely affects sampling efficiency and the accurate characterization of the posterior distribution.
Figure 10: Comparasion of three ABC-MCMC methods (ABC-MCMC, GL-ABC-MCMC with batch size 20 and global frequency $1$ and $0.2$). Error bars denote $95\%$ confidence intervals for $D_{KL}$ across 10 replicates. First row: $\epsilon$, $\mu$, $\sigma$; Second row: $\omega_{c}$, $\sigma_{c}$.
Figure 11: One-dimensional marginal distributions obtained by ABC-MCMC (first row), Global-Local ABC-MCMC with global frequency $\gamma=0.2$ (second row), and $\gamma=1$ (third row). The results correspond to 500,000 total iterations. The red line indicates the true value.
Figure 12: One-dimensional marginal distributions obtained by ABC-MCMC (first row), Global-Local ABC-MCMC with global frequency $\gamma=0.2$ (second row), and $\gamma=1$ (third row). The results correspond to the first 20,000 iterations of 500,000 total iterations. The red line indicates the true value.
@@ 7 Conclusion
In this article, we address the problem of Bayesian inference for models with intractable likelihood functions. Markov chain Monte Carlo (MCMC) methods are frequently combined with approximate Bayesian computation (ABC) to facilitate likelihood-free inference. However, MCMC algorithms relying solely on local moves often struggle to adequately explore posterior distributions with complex geometries. To overcome this limitation, we propose a novel framework that integrates both efficient global and local moves, thereby significantly enhancing the posterior exploration capability within the likelihood-free inference context.
Specifically, we develop a likelihood-free variant of the iterated Sequential Importance Resampling (i-SIR) algorithm to serve as an efficient global move within ABC-MCMC algorithms. For high-dimensional parameter spaces, we employ an ABC adaptation of the Metropolis-adjusted Langevin algorithm (MALA) as an effective local move, where the gradient of the log-likelihood function is approximated via the common random numbers technique.
To further improve the overall efficiency of the proposed ABC-MCMC scheme, we introduce two adaptive strategies. First, kernel density estimation is utilized to progressively refine the importance proposal distribution by introducing a sequence of intermediate target distributions. samsonov2022local-global proposed constructing the importance proposal distribution for i-SIR using normalizing flows. Although normalizing flows are powerful for modeling complex distributions, their performance can deteriorate when the initial proposal is far from the target distribution, particularly in ABC inference where many samples receive negligible weights, leading to unstable updates and computational inefficiency. Second, we define a criterion based on a unit cost version of the expected squared jumping distance (ESJD) to quantitatively assess the exploration capability of the ABC-MCMC kernel. Building on this, a sequential optimization algorithm is designed to select the hyperparameters governing the Global-Local ABC-MCMC moves.
Moreover, we rigorously establish the V-geometric ergodicity of the Global-Local ABC-MCMC algorithm that employs ABC-i-SIR as the global move. This theoretical result implies that, under suitable regularity conditions, the mixing rate of the proposed Markov kernel outperforms that of ABC-MCMC methods relying exclusively on local moves. Extensive numerical experiments on both synthetic and real-world models demonstrate that our algorithm achieves superior performance compared to existing approaches. A Python package glabcmcmc implementing this method is available at https://github.com/caofff/GL-ABC-MCMC. We also refer readers to cao2025glabcmcmc for more detailed information of this package.
Several avenues for future research and improvement arise from this work. First, the ABC tolerance parameter $\epsilon$ is currently pre-specified prior to running the ABC-MCMC algorithm. Following the framework proposed by del2012adaptive, an important extension is to integrate our Global-Local move scheme within an ABC sequential Monte Carlo (ABC-SMC) framework and explore adaptive strategies for tuning move parameters. Second, the computation of the Metropolis–Hastings acceptance probability, as well as the implementation of ABC-i-SIR, requires generating synthetic data from the simulation model, which can be computationally expensive for complex models. Future efforts will focus on accelerating ABC-MCMC inference by leveraging approximation techniques such as those in doi:10.1080/10618600.2024.2379349; Likelihood-freeapproximateGibbssampling. Third, the gradient estimation in approximate Bayesian methods for high-dimensional parameter spaces often requires a large number of simulations. To address this challenge, we plan to investigate variance reduction via control variates and develop methods to optimally determine simulation sizes for gradient approximation.
Acknowledge
This work was supported by the National Natural Science Foundation of China (12131001 and 12101333), the startup fund of ShanghaiTech University, the HPC Platform of ShanghaiTech University, the Fundamental Research Funds for the Central Universities, LPMC, and KLMDASR. The authorship is listed in alphabetic order.
SUPPLEMENTAL MATERIALS
Appendix A: